### v1

# Anton Alexa

Abstract
We establish a boundary delocalization principle for high–frequency Dirichlet eigenfunctions on smooth strictly convex domains. The main result is an exclusion of persistent boundary concentration at the level of individual eigenmodes relative to short spectral packets of sublinear length. Quantitatively, we compare boundary energies of single eigenfunctions to packet sums over frequency windows of size $N_{k}=o(k)$, without asserting any asymptotic gain in magnitude. The main mode–to–packet estimate relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain a packet zero–mean cancellation estimate. This mode–to–packet comparison is independent of eigenvalue monotonicity and is stable under eigenvalue crossings.
1 Introduction
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain, $d\geq 2$, and let $\{(\lambda_{k},u_{k})\}_{k\geq 1}$ denote the Dirichlet eigenpairs of $\Delta_{\Omega}$ with $\|u_{k}\|_{L^{2}(\Omega)}=1$. The boundary flux density $|\partial_{n}u_{k}|^{2}$ plays a central role in boundary spectral analysis and in microlocal restriction phenomena. A natural question is whether a single high–frequency eigenfunction can exhibit dominant boundary localization compared to its spectral neighbors.
Related results on boundary delocalization and restriction estimates typically rely on microlocal analysis, defect measures, or dynamical properties of the billiard flow, see for instance [5, 6, 12, 7, 1]. For general background on boundary eigenfunctions and related phenomena, see also the surveys [14, 15]. In contrast, the main mode–to–packet estimate is entirely static and relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain packet zero–mean cancellation.
This paper shows that, on strictly convex domains, persistent boundary concentration at the level of a single mode is excluded when one compares it to a spectral packet of sublinear length. We establish a packet–level delocalization principle: for any sublinear window $N_{k}=o(k)$, the ratio between a single boundary energy and the packet sum tends to zero. The proof uses the Rellich identity. In addition, we rule out collective boundary concentration: no positive fraction of modes inside a sublinear spectral packet can exhibit a persistent one–sided boundary bias with a fixed geometric weight of zero mean, in the spirit of averaged spectral quantities; see, e.g., [11, 10]. For the multi–mode statement we assume a slightly stronger growth condition on $N_{k}$ when $d\geq 3$.
Theorem 1.1 (Main result - Introductory Form).
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain, $d\geq 2$, and let $\{(\lambda_{k},u_{k})\}_{k\geq 1}$ be the Dirichlet eigenpairs of $\Delta_{\Omega}$. Let $N_{k}$ be any sublinear spectral window with $N_{k}\to\infty$ and $N_{k}=o(k)$. Then
$\frac{E_{k}}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\longrightarrow 0\quad\text{as }k\to\infty.$ (1)
In particular, no individual high–frequency eigenfunction can asymptotically carry a fixed positive proportion of the boundary energy of a sublinear spectral packet.
Remark 1.2 (Sharpness of the packet length condition).
The condition $N_{k}\to\infty$ is essential at the level of the present method. Indeed, for a fixed packet length $N$, the Rellich identity and the Weyl law alone do not imply that
$\frac{E_{k}}{\sum_{m=k}^{k+N-1}E_{m}}\longrightarrow 0.$ (2)
In particular, without averaging over a growing spectral window, a single eigenmode may contribute a non–negligible fraction of the packet boundary energy. Thus the requirement $N_{k}\to\infty$ represents the minimal averaging scale necessary to enforce mode–to–packet delocalization.
2 Preliminaries and notation
Let $\Omega\subset\mathbb{R}^{d}$ be bounded, $C^{\infty}$, and strictly convex. Denote by $\{(\lambda_{k},u_{k})\}_{k\geq 1}$ the Dirichlet eigenpairs of $\Delta_{\Omega}$, with $u_{k}|_{\partial\Omega}=0$ and $\|u_{k}\|_{L^{2}(\Omega)}=1$. For $x\in\partial\Omega$, define the boundary flux density
$\rho_{k}(x):=|\partial_{n}u_{k}(x)|^{2}.$ (3)
The total boundary energy is
$E_{k}:=\int_{\partial\Omega}\rho_{k}(x)\,d\sigma(x).$ (4)
Let $H$ denote the mean curvature of $\partial\Omega$ with respect to the outer normal and let $\bar{H}$ be its boundary average.
Definition 2.1 (Spectral packets).
Let $N_{k}$ be a positive integer sequence with $N_{k}\to\infty$ and $N_{k}=o(k)$. Define the index packet
$\mathcal{J}_{k}:=\{k,\,k+1,\,\dots,\,k+N_{k}-1\}$ (5)
and the packet boundary density
$\mathcal{Q}_{k,N_{k}}(x):=\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}\rho_{j}(x).$ (6)
For a measurable weight $w:\partial\Omega\to\mathbb{R}$, set
$E_{k}(w)$ $=\int_{\partial\Omega}\rho_{k}(x)\,w(x)\,d\sigma(x),$ (7)
$E_{k}^{abs}(w)$ $=\int_{\partial\Omega}\rho_{k}(x)\,|w(x)|\,d\sigma(x),$
$\mathcal{E}_{k,N_{k}}(w)$ $=\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}^{abs}(w).$
Definition 2.2 (Boundary correlation coefficient).
Let $w\in L^{\infty}(\partial\Omega)$ satisfy $\int_{\partial\Omega}w\,d\sigma=0$. Define
$C_{k}(w):=\frac{1}{E_{k}}\int_{\partial\Omega}\rho_{k}(x)\,w(x)\,d\sigma(x).$ (8)
For the multi–mode statement we assume $w\in C^{\infty}(\partial\Omega)$.
Remark 2.3 (Eigenvalue crossings).
All arguments are formulated at the level of boundary energy densities $\rho_{k}=|\partial_{n}u_{k}|^{2}$ and spectral sums. No ordering or differentiability of individual eigenvalues is required, and possible eigenvalue crossings do not affect the packet–level estimates.
3 Rellich identity and single–mode bounds
The Rellich identity yields the correct scaling for the boundary energy of a single eigenfunction on strictly convex domains; see, e.g., [13, 3].
Lemma 3.1 (Single–mode boundary energy bound).
There exist constants $0<c<C$ such that for all $k\geq 1$,
$c\,\lambda_{k}\leq E_{k}\leq C\,\lambda_{k}.$ (9)
Moreover, for any $w\in L^{\infty}(\partial\Omega)$,
$|E_{k}(w)|\leq E_{k}^{abs}(w)\leq C\,\lambda_{k}\,\|w\|_{L^{\infty}(\partial\Omega)}.$ (10)
Proof 3.2.
Fix $x_{0}\in\Omega$ and set $g(x):=\langle x-x_{0},\nu(x)\rangle$ on $\partial\Omega$. Since $\Omega$ is convex, for each $x\in\partial\Omega$ one has $\langle y-x,\nu(x)\rangle\leq 0$ for all $y\in\Omega$ (supporting hyperplane), where $\nu(x)$ is the outward unit normal. Since $x_{0}$ is interior, it cannot lie on the supporting hyperplane at $x$, hence $\langle x_{0}-x,\nu(x)\rangle<0$, and thus $g(x)=\langle x-x_{0},\nu(x)\rangle>0$ for all $x\in\partial\Omega$. By continuity and compactness, $0<m\leq g\leq M<\infty$ on $\partial\Omega$. The Rellich identity (see [13]) yields
$\int_{\partial\Omega}g(x)\,|\partial_{n}u_{k}(x)|^{2}\,d\sigma(x)=2\lambda_{k}.$ (11)
Therefore
$\frac{2}{M}\lambda_{k}\leq E_{k}\leq\frac{2}{m}\lambda_{k},$ (12)
which gives the claimed bounds. The weighted estimate follows from $|E_{k}(w)|\leq E_{k}^{abs}(w)\leq\|w\|_{L^{\infty}}E_{k}$. ∎
4 Packet scales and a contextual Weyl remark
The integrated boundary energy admits a Weyl-type leading asymptotic on smooth strictly convex domains; see, for instance, Safarov–Vassiliev and Ivrii. We record this only for context.
Definition 4.1 (Integrated boundary energy).
Define the partial boundary energy sum
$S(\Lambda):=\sum_{\lambda_{j}<\Lambda}E_{j}.$ (13)
Remark 4.2 (Integrated boundary Weyl expansion (context)).
It is known that
$S(\Lambda)=C_{\Omega}\,\Lambda^{1+\frac{d}{2}}+o\!\left(\Lambda^{1+\frac{d}{2}}\right),$ (14)
as $\Lambda\to\infty$; see, for example, [2, 4, 9]. This is recorded for context and is not used in the proof of the main theorem.
Lemma 4.3 (Packet zero–mean cancellation).
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain. Let $w\in C^{\infty}(\partial\Omega)$ satisfy $\int_{\partial\Omega}w\,d\sigma=0$. Let $N_{k}\to\infty$ with $N_{k}=o(k)$ and $N_{k}/k^{1-2/d}\to\infty$, and set $\mathcal{J}_{k}=\{k,\dots,k+N_{k}-1\}$. Then
$\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}(x)|^{2}\,w(x)\,d\sigma(x)=o\!\left(\sum_{j\in\mathcal{J}_{k}}E_{j}\right)\qquad(k\to\infty).$ (15)
Equivalently,
$\int_{\partial\Omega}\mathcal{Q}_{k,N_{k}}(x)\,w(x)\,d\sigma(x)=o\!\left(\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}\right)\qquad(k\to\infty).$ (16)
Proof 4.4.
For $\Lambda>0$ define the boundary spectral density
$Q_{\Lambda}(x):=\sum_{\lambda_{j}<\Lambda}|\partial_{n}u_{j}(x)|^{2},\qquad x\in\partial\Omega.$ (17)
By the boundary local Weyl law for the Dirichlet Laplacian paired with smooth test functions (see, e.g., [2, 4, 8]), there exist constants $A_{\Omega}>0$ and $B_{\Omega}\in\mathbb{R}$ such that
$\int_{\partial\Omega}Q_{\Lambda}(x)\,w(x)\,d\sigma(x)=A_{\Omega}\,\Lambda^{\frac{d+1}{2}}\int_{\partial\Omega}w\,d\sigma+B_{\Omega}\,\Lambda^{\frac{d}{2}}\int_{\partial\Omega}H\,w\,d\sigma+O\!\left(\Lambda^{\frac{d-1}{2}}\right),$ (18)
as $\Lambda\to\infty$, where $H$ is the mean curvature. Since $\int_{\partial\Omega}w\,d\sigma=0$, the leading term cancels. Hence there exists $C=C(\Omega,w)>0$ such that for all sufficiently large $\Lambda$,
$\left|\int_{\partial\Omega}Q_{\Lambda}(x)\,w(x)\,d\sigma(x)\right|\leq C\,\Lambda^{\frac{d}{2}}.$ (19)
Fix $k$ sufficiently large so that (19) applies to all $\Lambda\geq\lambda_{k-1}$. Choose $\Lambda$ with $\lambda_{k-1}<\Lambda<\lambda_{k}$. Then
$\int_{\partial\Omega}Q_{\Lambda}\,w\,d\sigma=\sum_{j<k}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma.$ (20)
Similarly, for $\Lambda^{\prime}$ with $\lambda_{k+N_{k}-1}<\Lambda^{\prime}<\lambda_{k+N_{k}}$,
$\int_{\partial\Omega}Q_{\Lambda^{\prime}}\,w\,d\sigma=\sum_{j<k+N_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma.$ (21)
Subtracting gives
$\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma=\int_{\partial\Omega}(Q_{\Lambda^{\prime}}-Q_{\Lambda})\,w\,d\sigma.$ (22)
By (19) and the triangle inequality,
$\left|\int_{\partial\Omega}(Q_{\Lambda^{\prime}}-Q_{\Lambda})\,w\,d\sigma\right|\leq\left|\int_{\partial\Omega}Q_{\Lambda^{\prime}}\,w\,d\sigma\right|+\left|\int_{\partial\Omega}Q_{\Lambda}\,w\,d\sigma\right|\leq C\Big((\Lambda^{\prime})^{\frac{d}{2}}+\Lambda^{\frac{d}{2}}\Big)\lesssim\lambda_{k+N_{k}}^{\frac{d}{2}}.$ (23)
On the other hand, by Lemma 3.1 and since $\lambda_{j}\geq\lambda_{k}$ for $j\in\mathcal{J}_{k}$,
$\sum_{j\in\mathcal{J}_{k}}E_{j}\geq c\sum_{j\in\mathcal{J}_{k}}\lambda_{j}\geq c\,N_{k}\,\lambda_{k}.$ (24)
Using Weyl asymptotics $\lambda_{k}\sim c_{\Omega}k^{2/d}$ and since $N_{k}=o(k)$, one has $\lambda_{k+N_{k}}/\lambda_{k}\to 1$. We obtain
$\frac{\left|\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma\right|}{\sum_{j\in\mathcal{J}_{k}}E_{j}}\lesssim\frac{\lambda_{k}^{\frac{d}{2}-1}}{N_{k}}\asymp\frac{k^{1-2/d}}{N_{k}},$ (25)
which tends to zero under the stated growth condition on $N_{k}$. This proves (15), and (16) follows by dividing by $N_{k}$. ∎
Remark 4.5 (Packet size condition).
For $d=2$ the condition $N_{k}/k^{1-2/d}\to\infty$ reduces to $N_{k}\to\infty$. For $d\geq 3$ it requires only mildly larger sublinear packets.
Remark 4.6 (Uniformity of constants).
All constants appearing in Lemma 3.1 and Proposition 4.7 are for the fixed strictly convex domain $\Omega$.
Proposition 4.7 (Packet average scale).
Let $N_{k}$ satisfy $N_{k}\to\infty$ and $N_{k}=o(k)$. Then there exist constants $c>0$ and $k_{0}$ such that for all $k\geq k_{0}$,
$\sum_{j\in\mathcal{J}_{k}}E_{j}\geq c\,N_{k}\,\lambda_{k}.$ (26)
Proof 4.8.
By Lemma 3.1 there exists $c_{0}>0$ such that $E_{j}\geq c_{0}\,\lambda_{j}$ for all $j$. Since $\lambda_{j}\geq\lambda_{k}$ for $j\in\mathcal{J}_{k}$, we obtain
$\sum_{j\in\mathcal{J}_{k}}E_{j}\geq c_{0}\,N_{k}\,\lambda_{k},$ (27)
which yields the lower bound. ∎
Remark 4.9 (Why sublinear packets).
The condition $N_{k}=o(k)$ means that the packet length is sublinear in the mode index and thus represents a short averaging window around $k$.
5 Boundary delocalization and spectral packet comparison
We now combine the packet scale lower bound with the single–mode upper bound.
Theorem 5.1 (Boundary delocalization via spectral packets).
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain. Let $N_{k}$ satisfy $N_{k}\to\infty$ and $N_{k}=o(k)$. Then
$\frac{E_{k}}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\longrightarrow 0\quad\text{as }k\to\infty.$ (28)
In particular, no single eigenmode can asymptotically carry a fixed positive proportion of the packet boundary energy on any sublinear spectral window.
Proof 5.2.
By Lemma 3.1, $E_{k}\leq C_{1}\lambda_{k}$. Proposition 4.7 gives $\sum_{m=k}^{k+N_{k}-1}E_{m}\geq c\,N_{k}\,\lambda_{k}$ for $k$ large. Hence
$\frac{E_{k}}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\leq\frac{C_{1}}{c\,N_{k}}\xrightarrow[k\to\infty]{}0,$ (29)
which yields the stated limit. ∎
Corollary 5.3 (Control of weighted boundary functionals).
Let $w\in L^{\infty}(\partial\Omega)$ be any bounded geometric weight. Then the weighted energy of a single mode is dominated by the unweighted packet sum:
$\frac{\left|\int_{\partial\Omega}|\partial_{n}u_{k}|^{2}\,w(x)\,d\sigma\right|}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\longrightarrow 0\quad\text{as }k\to\infty.$ (30)
Proof 5.4.
By Lemma 3.1, $|E_{k}(w)|\leq E_{k}^{abs}(w)\leq\|w\|_{\infty}E_{k}$. The result follows immediately from Theorem 5.1. ∎
Remark 5.5 (Weighted energies).
The same comparison can be extended to weighted boundary energies under additional assumptions on the weight; this is not used in the present paper. The geometric weight $w=|H-\bar{H}|$ is a natural example.
Theorem 5.6 (Absence of persistent one–sided boundary bias).
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain. Let $w\in C^{\infty}(\partial\Omega)$ satisfy
$\int_{\partial\Omega}w\,d\sigma=0.$ (31)
Let $N_{k}\to\infty$ with $N_{k}=o(k)$ and $N_{k}/k^{1-2/d}\to\infty$. Then for every $\varepsilon>0$,
$\frac{1}{N_{k}}\#\Big\{j\in\mathcal{J}_{k}:C_{j}(w)\geq\varepsilon\Big\}\longrightarrow 0\quad\text{as }k\to\infty.$ (32)
An analogous statement holds with $C_{j}(w)\leq-\varepsilon$.
Proof 5.7.
Suppose by contradiction that there exist $\varepsilon>0$, $\delta>0$, and a subsequence such that at least $\delta N_{k}$ indices in $\mathcal{J}_{k}$ satisfy $C_{j}(w)\geq\varepsilon$. Then
$\int_{\partial\Omega}\mathcal{Q}_{k,N_{k}}(x)\,w(x)\,d\sigma(x)=\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}\,C_{j}(w)\geq\varepsilon\delta\,\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}.$ (33)
By Proposition 4.7, the right–hand side is $\gtrsim\lambda_{k}$. By Lemma 4.3,
$\int_{\partial\Omega}\mathcal{Q}_{k,N_{k}}(x)\,w(x)\,d\sigma(x)=o\!\left(\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}\right).$ (34)
By Proposition 4.7 one has $\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}\gtrsim\lambda_{k}$, hence the left–hand side is $o(\lambda_{k})$. The case $C_{j}(w)\leq-\varepsilon$ is treated analogously. This contradicts the lower bound above and finishes the proof.
Corollary 5.8 (Vanishing density of strongly correlated modes).
For any fixed $w\in C^{\infty}(\partial\Omega)$ with zero mean, the set of indices $k$ such that
$\left|\int_{\partial\Omega}\rho_{k}(x)\,w(x)\,d\sigma(x)\right|\gtrsim E_{k}$ (35)
has asymptotic density zero.
Summary of the packet delocalization result
We have shown that on smooth strictly convex domains, single Dirichlet eigenfunctions cannot asymptotically dominate boundary energy relative to sublinear spectral packets. The main mode–to–packet estimate is purely static and relies only on the Rellich identity.
Conclusion
We have proved a quantitative boundary delocalization principle for Dirichlet eigenfunctions on smooth strictly convex domains, showing that no individual high–frequency eigenmode can asymptotically dominate the boundary energy of a sublinear spectral packet. The result excludes persistent boundary concentration at the level of single eigenfunctions when compared to neighboring modes in vanishing relative frequency windows.
The main mode–to–packet estimate is entirely static and relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law; integrated boundary Weyl asymptotics for the Dirichlet Laplacian [2, 4, 8] are recorded for context. In particular, no microlocal defect measures, dynamical assumptions on the billiard flow, or ergodicity hypotheses are required. The argument is robust under eigenvalue crossings and depends only on global spectral information.
From a conceptual viewpoint, the result provides a mode–to–packet comparison mechanism on the boundary, analogous to bulk delocalization phenomena but valid in a purely deterministic and geometric setting. The sublinear packet scale reflects the minimal averaging required to control single–mode contributions using Rellich-type bounds. In addition, we obtain a multi–mode delocalization principle: no positive fraction of modes in a sublinear packet can maintain strong boundary correlation with a fixed zero–mean weight.
Several directions for further investigation suggest themselves. First, it would be natural to extend the argument to other boundary conditions or to more general elliptic operators with boundary. Second, one may ask whether analogous packet–level delocalization phenomena hold on nonconvex domains or on Riemannian manifolds with boundary, where boundary Weyl remainders are less rigid. Finally, the present comparison principle may serve as a useful input in the study of boundary spectral statistics and correlations, complementing existing microlocal and dynamical approaches.
We expect that the mode–to–packet perspective introduced here can be applied in a broader range of boundary spectral problems where individual eigenfunction control is too weak, but averaged spectral information remains accessible.